A can containing Freon gas (CCl2F2) at 2.14 atm and 12 ∘C is thrown into a fire which is at 740. ∘C. What will be its pressure in atmospheres at 740. ∘C?

Express your answer to three significant figures.
Answer #1

According to Gas law,

P1/T1 = P2/T2

P1 & P2 = pressures at temperature T1 and T2.

P1 = 2.14 atm. T1 = 12°C

P2 = ? T2 = 740°C

2.14/12 = P2/740

P2 = (2.14×740)/12

P2 = 132 atm

Answer = 132 atm